What is the solution of - .744 ? e) 2 d) 4 b) 0 a) -2
Answer #1

Correct answer is a

X xty 6) = 7 x+y a - 7 xty 1 = 7*7 72 x x txty 1 = 1 = 72x+y → 7=72274 7 A =) 0= 2x+y 2212 - ។ -> x= -4 2 x = -2